# Effect of freestream turbulence on the coherent dynamics of a wind turbine wake

### v1

@@ Abstract
The wake of a model wind turbine exposed to freestream turbulence (FST) with a variety of turbulent characteristics is studied through Particle Image Velocimetry experiments. The FST cases spanned turbulent intensities ($T_{i}$) in the range $1\%\lesssim T_{i}\lesssim 13.1\%$ and integral length scales ($L_{v}$) in the range $0.01D\lesssim L_{v}\lesssim 0.11D$ (where $D$ is the diameter of the turbine model). Increasing $T_{i}$ and $L_{v}$ in this range resulted in an earlier breakdown of the tip vortices which in turn resulted in an earlier onset of wake recovery. Typically, the amplitude of wake meandering increased in the presence of FST, but not monotonically with either $T_{i}$ or $L_{v}$. The nature of the low frequency region (typical of wake meandering frequencies) of the power spectra changed from a more discrete to a more broadband one as FST level was increased, highlighting a fundamental change in the nature of wake meandering with/without FST. Deeper insights into the origins, and subsequent evolution, of the various coherent motions (characterised by particular frequencies) in the presence of FST are obtained through analysis of the multi-scale triple-decomposed coherent kinetic energy budgets. The wake meandering modes in the presence of FST are shown to better utilize the mean velocity shear, extracting more energy from the mean flow while other sources such as non-linear triadic interactions and diffusion also gain in importance.
@@ 1 Introduction
The effect of FST has been studied through a combination of numerical simulations (largely using large eddy simulation, LES), laboratory experiments using a small-scale representative wind turbine, and field experiments. High levels of FST has been shown to break down the tip vortices in the near wake [Chamorro and Porté-Agel, 2009, Wu and Porté-Agel, 2012]. These vortices act as a ‘shield’ in the near wake, inhibiting the exchange of mass, momentum and kinetic energy [Medici, 2005]. Lignarolo et al. [2015] performed particle image velocimetry (PIV) experiments on a two-bladed wind turbine model and showed that there was a net kinetic energy entrainment into the wake only after the tip vortices broke down. The recent study by Biswas and Buxton [2024b] reached a similar conclusion for a three-bladed wind turbine model. Therefore, high FST levels can accelerate wake recovery by breaking down this shield of vortices. In fact, a linear relationship between $T_{i}$ and the wake recovery rate has been reported [Niayifar and Porté-Agel, 2016, Carbajo Fuertes et al., 2018, Gambuzza and Ganapathisubramani, 2023].
@@
In the near wake, freestream turbulence breaks down the dominant coherent structures i.e. the tip vortices. In contrast, freestream turbulence likely plays a different role on the dominant far wake structures, which have long been primarily associated with wake meandering. Although the dominance of wake meandering in the far wake has been known for many years, there still remain varied opinions on the genesis of the meandering motion. For instance, a distinct wake meandering frequency has been reported when freestream turbulence intensity was negligible [Okulov et al., 2014, Biswas and Buxton, 2024a] thereby associating wake meandering to a rotor scale bluff body shedding type instability [Medici and Alfredsson, 2008, Howard et al., 2015, Biswas and Buxton, 2024a]. In contrast, in other works wake meandering has been treated as a ‘random’ unsteady oscillation of the far wake driven by freestream turbulence [Porté-Agel et al., 2020]. Here, the wind turbine wake is thought to be passively advected by the large scale structures present in the freestream turbulence [Larsen et al., 2008]. Unifying the two schools of thoughts, there is now a growing evidence that in real wind turbines exposed to FST, the nature and intensity of wake meandering depends on multiple factors including the turbine geometry, operating condition and freestream conditions [Howard et al., 2015, Mao and Sørensen, 2018, Heisel et al., 2018, Gupta and Wan, 2019, Gambuzza and Ganapathisubramani, 2023, Biswas and Buxton, 2024a].
Although, the effect of $T_{i}$ on a turbine wake has been studied extensively, the need to characterize the spectral content of the inflow turbulence has been realized only recently [Mao and Sørensen, 2018, Gupta and Wan, 2019, Vinod and Banerjee, 2019, Gambuzza and Ganapathisubramani, 2023, Hodgson et al., 2023, Bourhis et al., 2025]. In engineering wake models, the growth of the wake is often assumed to be a function of turbulence intensity alone and the spectral characteristic of the incoming flow is often ignored [Gambuzza and Ganapathisubramani, 2023]. Low frequencies present in the inflow have been shown to modulate the instantaneous power output of a stand alone turbine [Chamorro et al., 2013]. There is a newly established consensus that wind turbines do not simply advect the structures present in the freestream as was originally assumed by the dynamic wake meandering model [Larsen et al., 2008]. Instead, a turbine has a dynamic response i.e. certain frequencies in the freestream can be amplified in the wake [Mao and Sørensen, 2018, Heisel et al., 2018, Gupta and Wan, 2019, Hodgson et al., 2023]. Mao and Sørensen [2018] identified nonlinear optimal perturbations to a wind turbine wake modeled as an actuator disk. They reported that as the inflow perturbation amplitude increased, the optimal perturbation frequency reduced from $St=0.64$ to $St=0.25$ for $Re_{D}=1000$. As $Re_{D}$ was increased to 3000, the optimal perturbation frequency further reduced to $St=0.16$. These low frequencies fall well within the accepted range of wake meandering frequencies [Medici and Alfredsson, 2008, Chamorro et al., 2013, Okulov et al., 2014]. Beyond $Re=3000$, Mao and Sørensen [2018] noted that the optimal frequency did not change, indicating a possible $Re$ invariance at all application specific Reynolds numbers. Similarly, through field experiments, Heisel et al. [2018] reported that the wind turbine amplifies a range of frequencies in the inflow yielding spectral peaks in the range $St=0.3-0.4$ in the far wake. Similarly, larger freestream characteristic length scales were shown to induce a stronger wake meandering and were linked to a faster kinetic energy entrainment from the freestream [Blackmore et al., 2014, Vahidi and Porté-Agel, 2024]. However, these studies modeled the turbine as an actuator disk so, the effect of the tip vortices was not considered. The results of these studies likely hold better in the far wake of a turbine which is better approximated by the wake of an actuator disk. Interestingly, Hodgson et al. [2023] reported a shorter near wake region and faster wake recovery for shorter integral time scales. These results indicate that the freestream scales may have a different effect in the near and the far wake [Hodgson et al., 2023]. The current study aims to further understand the effect of the freestream characteristic length scales on the near wake of a turbine.
@@
Unlike several previous works where the statistical properties of the wake were studied in the presence of FST, we primarily focus on the dynamic properties. A series of time-resolved PIV experiments are performed on a wind turbine model exposed to freestream turbulence having different turbulence intensities ($T_{i}$) and integral length scales ($L_{v}$). In this work we build on our previous works by including freestream turbulence and we attempt to better understand the following aspects: (a) How does the energy content at the dominant frequencies in the near-wake change with freestream turbulence? (b) In the absence of freestream turbulence, Biswas and Buxton [2024a] found a scaling between the wake meandering frequency and $\lambda$ indicating that wake meandering could be a rotor scale bluff body shedding type instability. How does the nature of wake meandering change in the presence of freestream turbulence? (c) How does freestream turbulence affect the energy exchanges between the various coherent modes? (d) Finally, in all of the above what are the relative roles of $T_{i}$ and $L_{v}$?
Figure 1: (a) Parameter space of turbulence intensity ($T_{i}$) and integral length scale ($L_{v}$) produced by different turbulence generating grids. The error bars show the maximum and minimum values observed within the rotor diameter. (b) shows the effective tip speed ratios ($\lambda_{eff}$) for $\lambda_{\infty}=6$. $\lambda_{eff}$ is lower than $\lambda_{\infty}$ as the freestream velocity increases due to the blockage caused by the turbulence generating grids. The uncertainties in $\lambda_{eff}$ are due to the variation of the freestream velocity within the rotor diameter.
@@ 2 Methodology
Particle image velocimetry (PIV) experiments were conducted in the hydrodynamics flume in the Department of Aeronautics at Imperial College London. The flume had a cross-sectional area of $60\times 60$ cm2 at the operating water depth. The turbine radius ($R$) was 0.1 m and the model was the same as that detailed in [Biswas and Buxton, 2024a, b]. The model had a nacelle and tower associated with it such that it resembled a utility-scale turbine. The free-stream velocity ($U_{\infty}$) was kept constant at 0.2 ms-1. Experiments were performed at a single tip speed ratio $\lambda=6$. The Reynolds number based on the turbine diameter $Re_{D}$ (where $D$ is the turbine diameter) was $\approx 40000$.
Different freestream turbulence cases were obtained by using different passive turbulence generating grids similar to that used in our previous studies [Kankanwadi and Buxton, 2020, Chen and Buxton, 2023]. Separate PIV experiments were conducted to characterise the variation of turbulence intensity ($T_{i}$), integral length scales ($L_{v}$) and the mean streamwise velocity as a function of streamwise distance from the grids. The details of these experiments can be found in appendix 1. The parametric space (spanned by $T_{i}$ and $L_{v}$) in this study is shown in figure 1(a). Specific symbols are used for certain types of grids. The symbols represent the value of $T_{i}$ and $L_{v}$ at hub height, while the error bars show the maximum and minimum values observed within the rotor diameter (see the transverse variations of $T_{i}$ and $L_{v}$ in figures 23 and 25 respectively in appendix 1).
@@
The grids also introduce a blockage to the freestream, resulting in an acceleration of the incoming flow. Although the tip speed ratio based on $U_{\infty}$ (denoted as $\lambda_{\infty}$) was kept constant at 6 for all the FST cases, due to the blockage induced increment of the freestream velocity, the effective tip speed ratio ($\lambda_{eff}$) reduced for all the FST cases. $\lambda_{eff}$ for the different cases is shown in figure 1(b). The symbols represent the values at the geometrical centreline whilst, the uncertainties are due to the transverse variation of the mean streamwise velocity. The uncertainties increase with $T_{i}$ as can be expected. The test cases are divided into 4 subgroups, primarily based on turbulence intensity. Group 1 has negligible $T_{i}$ while groups 2, 3 and 4 have low, mild and high $T_{i}$ respectively. This classification was done by comparing $T_{i}$ as well as the nature of the tip vortex related frequencies observed in the near wake as will be discussed in the upcoming sections. Nevertheless, hard boundaries likely do not exist between the adjacent subgroups.
Figure 2: Exemplar instantaneous vorticity fields for the cases (a) 1, (b) 2a, (c) 3, and (d) 4b.
@@
The experiments with the wind turbine had the same field of view as that in experiment 1A discussed in [Biswas and Buxton, 2024b]. Therefore, the details of the experiments are only briefly discussed here. Only planar PIV experiments were performed at the diametrical plane of the rotor aligned with the tower’s axis. The experiments were time resolved. The acquisition frequency ($100$ Hz) was much higher than the turbine’s rotational frequency ($\approx 1.87$ Hz). The spatial resolution of the experiments ($\approx 0.0115D$) was good enough to capture the tip vortices (having size $\approx 0.07D$).
3 Vorticity field
Instantaneous vorticity fields for 4 cases, one from each subgroup (increasing turbulence intensity) are shown in figure 2 for $\lambda_{\infty}=6$ (see also supplementary videos 1-4). For case 1, the tip vortices undergo merging at $x/D\approx 1.5$ similarly to the observation in our earlier study (in the absence of FST) through a different set of experiments [Biswas and Buxton, 2024a]. Due to the outer fluid being non-turbulent, the wake edge is clearly marked by vorticity in case 1. The solid black line shows the mean wake centreline $y=y_{c}$ obtained by averaging over instantaneous wake centrelines (the loci of the minima in the instantaneous streamwise velocity profiles). The centreline bends towards the $y<0$ side due to the presence of the tower as was also discussed in previous studies [Pierella and Sætran, 2017, Biswas and Buxton, 2024a]. For low $T_{i}$ (case 2), tip vortices and their merging is still observed, however the trajectories of both the individual and merged vortices appear more erratic similar to the observation of Gambuzza and Ganapathisubramani [2023]. Large scale structures from the freestream penetrate deep into the wake, close to the wake centreline (see the supplementary video 2). As $T_{i}$ is increased further, the tip vortices break down much closer to the turbine as reported in previous studies [Chamorro and Porté-Agel, 2009, Wu and Porté-Agel, 2012, Gambuzza and Ganapathisubramani, 2023]. The presence of background turbulence makes it difficult to visually identify the wake edge for both medium and high $T_{i}$ cases.
Figure 3: Spectra at $x/D=0.5$ and $y/D=0.55$ for the cases (a) 1, (b) 2a, (c) 3, and (d) 4b.
4 Effect of freestream turbulence on the important frequencies

### v2

@@ Abstract
The wake of a model wind turbine exposed to incoming freestream turbulence (FST) with a variety of turbulent characteristics is studied through Particle Image Velocimetry experiments. The FST cases were produced using different passive turbulence generating grids. The cases spanned turbulent intensities ($T_{i}$) in the range $1.3\%\lesssim T_{i}\lesssim 14\%$ and only considered short integral length scales $L_{v}\lesssim 0.2D$ (where $D$ is the turbine diameter). Increasing $T_{i}$ and $L_{v}$ in this range resulted in an earlier breakdown of the tip vortices which in turn resulted in an earlier onset of wake recovery. For all the FST cases considered, the initiation of wake meandering was found to be related to an intrinsic instability of the turbine, even for the cases with the highest FST levels. The amplitudes of wake meandering were similar for all the cases in the near wake ($x<2D$), but the amplitudes in the far wake ($x>4D$) were discernibly higher for all the FST cases compared to the no grid case (lowest $T_{i}$), primarily due to the early break down of the tip vortices. Deeper insights into the origins, and subsequent evolution, of the various coherent motions (characterised by particular frequencies) in the presence of FST are obtained through analysis of the multi-scale triple-decomposed coherent kinetic energy budgets. The wake meandering modes in the presence of FST are shown to better utilize the mean velocity shear, extracting more energy from the mean flow while other sources such as non-linear triadic interactions and diffusion also become important.
@@ 1 Introduction
The effect of FST has been studied through a combination of numerical simulations (largely using large eddy simulation, LES), laboratory experiments using a small-scale representative wind turbine, and field experiments. High levels of FST have been shown to break down the tip vortices in the near wake [Chamorro and Porté-Agel, 2009, Wu and Porté-Agel, 2012]. These vortices act as a ‘shield’ in the near wake, inhibiting the exchange of mass, momentum and kinetic energy [Medici, 2005]. Lignarolo et al. [2015] performed particle image velocimetry (PIV) experiments on a two-bladed wind turbine model and showed that there was a net kinetic energy entrainment into the wake only after the tip vortices broke down. Similar conclusions were reached for a three-bladed wind turbine [Eriksen and Krogstad, 2017, Biswas and Buxton, 2024b]. Therefore, high FST levels can accelerate wake recovery by breaking down this shield of vortices. In fact, a linear relationship between $T_{i}$ and the wake recovery rate has been reported [Niayifar and Porté-Agel, 2016, Carbajo Fuertes et al., 2018, Gambuzza and Ganapathisubramani, 2023].
@@
In the near wake, freestream turbulence breaks down the dominant coherent structures i.e. the tip vortices. In contrast, freestream turbulence likely plays a different role on the dominant far wake structures, which have long been primarily associated with wake meandering. Although the dominance of wake meandering in the far wake has been known for many years, there still remain varied opinions on the genesis and characteristics of the meandering motion. For instance, a distinct wake meandering frequency has been reported when freestream turbulence intensity was negligible [Okulov et al., 2014, Biswas and Buxton, 2024a] thereby associating wake meandering to a rotor scale bluff body shedding type instability [Medici and Alfredsson, 2008, Howard et al., 2015]. In contrast, in other works wake meandering has been treated as a ‘random’ unsteady oscillation of the far wake driven by freestream turbulence [Porté-Agel et al., 2020]. Here, the wind turbine wake was thought to be passively advected by the large scale structures present in the freestream turbulence [Larsen et al., 2008]. However, later works suggested that wind turbines do not simply advect the structures present in the freestream, instead, a turbine has a dynamic response i.e. certain frequencies in the freestream can be amplified in the wake [Chamorro et al., 2012, Singh et al., 2014, Mao and Sørensen, 2018, Heisel et al., 2018, Gupta and Wan, 2019, Hodgson et al., 2023], highlighting the importance of characterising the spectral content of the inflow. Chamorro et al. [2012] proposed that a turbine can be modeled as an ‘active filter’, generating or amplifying certain frequencies while dampening others. Similar results were obtained by Singh et al. [2014] who showed that the turbine made the velocity fluctuations in the wake more ‘homogenised’ compared to the incoming flow.
Mao and Sørensen [2018] identified nonlinear optimal perturbations to a wind turbine wake modeled as an actuator disk. They reported that as the inflow perturbation amplitude increased, the optimal perturbation frequency reduced from $St=0.64$ to $St=0.25$ (where $S_{t}$ denotes the Strouhal number based on the turbine diameter and freestream velocity) for $Re=1000$ (based on turbine radius and freestream velocity). As $Re$ was increased to 3000, the optimal perturbation frequency further reduced to $St=0.16$. These low frequencies fall well within the accepted range of wake meandering frequencies [Medici and Alfredsson, 2008, Chamorro et al., 2013, Okulov et al., 2014]. Beyond $Re=3000$, Mao and Sørensen [2018] noted that the optimal frequency did not change, indicating a possible $Re$ invariance at all application specific Reynolds numbers. Similarly, the LES study of Foti et al. [2018] showed that the wake meandering amplitude, when properly scaled by the thrust, was largely invariant of the rotor size that spanned nearly 3 orders of magnitude (see figure 8(b) of Foti et al. [2018]). The authors however, treated wake meandering as more of a bluff body shedding type instability of the turbine [Howard et al., 2015]. Therefore, there is now a general consensus that in real wind turbines exposed to FST, the nature and intensity of wake meandering can depend on multiple factors including the turbine geometry, operating condition and freestream turbulence [Howard et al., 2015, Mao and Sørensen, 2018, Gupta and Wan, 2019, Gambuzza and Ganapathisubramani, 2023, Dong et al., 2023, Biswas and Buxton, 2024a, Wang et al., 2025]. The relative importance of each of these factors is still not fully understand and can well differ in the near wake and the far wake of the turbine [Foti et al., 2018, Dong et al., 2023, Bourhis et al., 2025]. Dong et al. [2023] studied three different blade designs under three different FST conditions but at similar $C_{T}$. Within $x<3D$, the wake meandering amplitudes (measured using the standard deviation of the wake centreline) were similar for the different blade designs and turbulence levels. Further downstream however, the blade design producing the strongest hub vortex showed significantly stronger wake meandering. For all the blade designs, in the far wake, the high FST case showed slightly higher wake meandering compared to low/medium FST cases (see figure 19 of [Dong et al., 2023]).
Figure 1: Schematic of the experimental setup.
While the need to characterise the spectral content of the inflow has been realised for some time, the exact role of the freestream length/time scales on the dynamics of the wind turbine wake and the wake recovery process is still not fully understood. Larger freestream characteristic length scales were shown to induce a stronger wake meandering and were linked to a faster kinetic energy entrainment from the freestream towards the far wake [Blackmore et al., 2014, Vahidi and Porté-Agel, 2024]. These studies however modeled the turbine as an actuator disk, hence the effect of the tip vortices in the near field was ignored. Interestingly, Hodgson et al. [2023] reported a shorter near wake region and faster wake recovery in wind turbines exposed to a flow with shorter integral time scales. The authors modeled the wind turbine using the actuator line method (i.e. the tip vortices were captured). Similarly, the PIV experiments by Gambuzza and Ganapathisubramani [2023] showed that for very large integral time scales in a non-Kolmogorov type flow, the wake recovered slowly compared to an inflow with a similar $T_{i}$ but shorter integral time scale. The authors linked this observation to the delayed break-up of the tip shear layer for this type of flows. Interestingly however, the recent experimental work by Bourhis et al. [2025] did not find any clear influence of the freestream integral length scales on the strength of the leading tip vortex frequencies. These results indicate that the effect of the freestream characteristic length scales on the near and far wake dynamics of a turbine and the wake recovery process is still far from clear.
@@
In this work, we explore answers to some of these appealing questions by performing a series of time-resolved PIV experiments on a wind turbine model exposed to incoming freestream turbulence having different turbulence intensities ($T_{i}$) and integral length scales ($L_{v}$). The wind turbine model included a nacelle and tower mimicking a real wind turbine. In particular, we attempt to better understand the following aspects: (a) How does the energy content at the dominant frequencies in the near-wake change with freestream turbulence? (b) In the absence of freestream turbulence, Biswas and Buxton [2024a] found a scaling between the wake meandering frequency and $\lambda$ indicating that wake meandering could be a rotor scale bluff body shedding type instability. How does the nature of wake meandering change in the presence of freestream turbulence? (c) How does freestream turbulence affect the energy exchanges between the various coherent modes and how does it impact the wake recovery process? (d) Finally, in all of the above what are the relative roles of $T_{i}$ and $L_{v}$?
@@ 2 Methodology
The PIV experiments with the wind turbine had the same field of view as that in experiment 1A discussed in [Biswas and Buxton, 2024b]. Therefore, the details of the experiments are only briefly discussed here. Only planar PIV experiments were performed at the diametrical plane of the rotor aligned with the tower’s axis. Three phantom v641 cameras were used to yield a stitched field of view (FOV) spanning $\approx 4.5D$ in the streamwise ($x$) direction and $\approx 1.1D$ in the spanwise ($y$) direction. The origin of the co-ordinate system was placed at the rotor plane at the hub height. The experiments were time resolved. The acquisition frequency ($100$ Hz) was much higher than the turbine’s rotational frequency ($\approx 1.87$ Hz). The spatial resolution of the experiments ($\approx 0.0115D$) was good enough to capture the tip vortices (having size $\approx 0.07D$). Data was obtained for a total time of 54.56 seconds (5457 snapshots) which corresponded to $\approx 102$ rotor revolutions.
Figure 2: Profiles of mean streamwise velocity at the rotor plane (with the turbine removed). The case IDs are shown on the top right.
Figure 3: Turbulent intensity profiles at the rotor plane (with the turbine removed). The case IDs are shown on the bottom right.
Figure 4: Profiles of integral length scales at the rotor plane (with the turbine removed). The case IDs are shown on the bottom right.
In previous experimental studies, different ‘flavours’ of freestream turbulence, primarily characterised by different turbulence intensities ($T_{i}$) and freestream integral length scales ($L_{v}$), has been produced by either using a passive turbulence generating grid [Blackmore et al., 2016, de Oliveira et al., 2025] or using an active turbulence generating grid [Gambuzza and Ganapathisubramani, 2023, Bourhis et al., 2025]. Active grids have been shown to be capable of producing turbulent Reynolds numbers and integral length scales much larger than those possible using passive grids [Kurian and Fransson, 2009, Hearst and Lavoie, 2015, Hearst et al., 2018]. Despite limitations, passive grids have been extensively used in the literature primarily due to the ease of manufacturing and installation [Bearman and Morel, 1983, Fransson et al., 2005, Kankanwadi and Buxton, 2020]. Using both types of grids however, it is generally difficult to fill the top-left and bottom-right corners of a $T_{i}$ - $L_{v}$ map, i.e. FST with very large integral length scales and very small turbulent intensities and vice versa.
Figure 5: (a) Parameter space of turbulence intensity ($T_{i}$) and integral length scale ($L_{v}$) produced by different turbulence generating grids. The symbols show the values averaged over the rotor radius. (b) shows the effective tip speed ratios ($\lambda_{eff}$) for $\lambda_{\infty}=6$. $\lambda_{eff}$ is measured based on the rotor averaged bulk velocity, $U_{b}$ instead of $U_{\infty}$ and hence is lower than $\lambda_{\infty}$. The same symbol is used for a particular type of grid used as indicated in (a). The case IDs are also indicated in (a).
[…]
Table 1: Parameters associated with the different experiments. Here $X_{G}$ denotes the distance between the grid and the turbine. $x^{*}$ is the wake intersection length scale for the fractal grids [Mazellier and Vassilicos, 2010] and $M$ is the mesh size of the regular grids. $U_{b}$ denotes the rotor averaged bulk velocity. $\lambda_{eff}$ is the tip speed ratio based on $U_{b}$.
[…]
In table 1, the distance between the grid and the turbine ($X_{G}$) is also shown in terms mesh size ($M$) for the regular grids and the wake intersection length scale ($x^{*}$) for the fractal grids. Here $x^{*}$ is defined as $x^{*}=L_{0}^{2}/t_{0}$, where $L_{0}$ and $t_{0}$ represent the length and thickness of the largest bars of the fractal grids [Mazellier and Vassilicos, 2010, Kankanwadi and Buxton, 2020]. For the regular grids there is a monotonic decay of turbulence and the turbulence becomes more homogeneous with downstream distance from the grid. For the fractal grids on the other hand, turbulence develops much slowly. For the latter, $T_{i}$ generally peaks around $x=0.45x^{*}$ and the flow starts to become homogeneous at around $x=0.5x^{*}$ [Mazellier and Vassilicos, 2010]. Considering several additional experimental constraints however, the grids had to be placed at locations with respect to the turbine where the inflow turbulence is not homogeneous, especially for the highest FST cases. Additionally for some of the cases, especially cases 2c and 4c (closer to the regular grids) $T_{i}$, $L_{v}$ and $\bar{U}$ also had a considerable streamwise gradient. The streamwise variation of $T_{i}$, $L_{v}$ and $\bar{U}$ behind all the turbulence generating grids is shown in the supplementary material. Nevertheless, the inhomogeneity of the inflow turbulence likely only has a secondary role on the dynamics of the wake as we will be discussing.
One of the other factors that had to be considered while positioning the grid was the interaction of the grid with the boundary layer developing at the bottom surface of the flume. For instance, the regular grid SRG38 was found to strongly interact with the boundary layer. This was most likely because this grid had horizontal bars closest to the bottom wall of the flume (highest local blockage) in the central laser plane (see figure 1 in the supplementary material). It can be seen from figure 2, how the boundary layer gets progressively thicker as we move downstream of the grid (from figure 2(d) - 2(b)). Therefore, further away from the grid, although we could gain in terms of homogeneity, the boundary layer would introduce unwanted additional turbulence. For the case 2a, although the turbine was placed outside the boundary layer at the rotor plane, further downstream ($x>3D$), the turbine wake intersected the boundary layer. Therefore, for this particular case, the results were only discussed up to $x/D=3$. This was less of an issue for the other grids. The regular grid SRG111 was intentionally placed closer to the turbine to create the cases 3b and 4c with relatively higher $T_{i}$ but at the expense of homogeneity. For the fractal grids on the other hand, since the turbulence develops much more slowly, a much higher $X_{G}$ was necessary. Now, the turbine had to be placed towards the upstream end of the test section in order to have a large portion of the wake captured in the test section measurement area (see figure 1). This resulted in the distance between the turbine and the upstream end of the flume being limited to $\approx$ 3m ($15D$). So the highest $X_{G}$ tested without shortening the streamwise extent of the wake measured was $15D$. Only for experiment 4a, the fractal grid SFG7 was placed at $X_{G}=17D$, but with the wake measured up to $x/D=3$ to see if there is a significant change in the near wake dynamics.
Figure 6: Exemplar instantaneous vorticity fields for the cases (a) 1, (b) 2a, (c) 2c, (d) 3b, (e) 4b and (f) 4d. The black dash-dotted line highlights the wake edge towards the far wake.
Estimating integral length scales involves inherent intricacies and different methods yield nonidentical results in an inhomogeneous anisotropic turbulent flow as in the present case. For instance, the integral length scales can be obtained by the direct computation of the spatial auto-correlation function or by obtaining integral time scales using the temporal auto-correlation and then invoking Taylor’s hypothesis of frozen turbulence. In this work, we follow the former approach. The integral length scales are estimated as $L_{v}=\int_{0}^{\hat{r}}R_{12}^{{}^{\prime}}dr$, where $R_{12}^{{}^{\prime}}$ is the transverse autocorrelation function of the streamwise velocity fluctuations. $R_{12}^{{}^{\prime}}$ is obtained using the following equation:
$R_{12}^{{}^{\prime}}(\textbf{x},\textbf{y},r)=\frac{\overline{u_{1}^{\prime}(\textbf{x},\textbf{y})u_{1}^{\prime}(\textbf{x},\textbf{y}+r\textbf{e}_{2})}}{\sqrt{\overline{u_{1}^{\prime}(\textbf{x},\textbf{y})^{2}}}\sqrt{\overline{u_{1}^{\prime}(\textbf{x},\textbf{y}+r\textbf{e}_{2})^{2}}}}$ (1)
Here the overbars denote averaging over time. $u_{1}^{\prime}$ is the streamwise velocity fluctuation and $r\textbf{e}_{2}$ represents some displacement $r$ along the $\textbf{e}_{2}$ direction ($\textbf{e}_{2}$ is the unit vector along the spanwise, $y$ direction). Finally $L_{v}$ is calculated by integrating $R_{12}^{{}^{\prime}}$ up to a certain distance $\hat{r}$ where $R_{12}^{{}^{\prime}}$ first becomes zero. Further details about the quantification of the integral length scales and the sources of uncertainties have been added in appendix 1 and in the supplementary material. It is worthwhile mentioning that other transverse or streamwise autocorrelation functions were also tested. It was seen that the streamwise autocorrelation function of the transverse velocity ($R_{21}^{{}^{\prime}}$) yielded integral length scales of magnitude similar to that obtained using $R_{12}^{{}^{\prime}}$ whilst, the streamwise autocorrelation function of the streamwise velocity ($R_{11}^{{}^{\prime}}$) yielded larger (up to 232-3 times) integral length scales. However the autocorrelation functions $R_{12}^{{}^{\prime}}$ or $R_{21}^{{}^{\prime}}$ more often fell off to zero compared to $R_{11}^{{}^{\prime}}$ which exhibited long range correlations. Therefore, only the integral length scales calculated using $R_{12}^{{}^{\prime}}$ are discussed here. A more detailed discussion on the different correlations can be found in the supplementary material. Nevertheless, the largest integral length scales we are able to produce here (<0.2D<0.2D using $R_{12}^{{}^{\prime}}$ or <0.45D<0.45D using $R_{11}^{{}^{\prime}}$) are an order of magnitude smaller than the largest length scales a real wind turbine residing in an atmospheric boundary layer can be exposed to [Larsen et al., 2007, Porté-Agel et al., 2020]. The current study is thus only able to explore the effect of short FST length scales on the near wake dynamics of a turbine.
The grids also introduce a blockage to the freestream, resulting in an acceleration of the incoming flow (please see the supplementary material). Although the tip speed ratio based on $U_{\infty}$ (denoted as $\lambda_{\infty}$) was kept constant at 6 for all the FST cases, due to the blockage induced increment of the freestream velocity (see figure 2), the effective tip speed ratio (defined as $\lambda_{eff}=\lambda_{\infty}U_{\infty}/U_{b}$) reduced for all the FST cases. Here $U_{b}$ is the rotor averaged bulk velocity. $\lambda_{eff}$ for the different cases is shown in figure 5(b). The uncertainties are due to the transverse variation of the mean streamwise velocity as shown in figure 2. The variation of the power ($C_{P}$) and thrust coefficients ($C_{T}$) as a function of the tip speed ratio were obtained in our previous work using the blade element momentum method [Biswas and Buxton, 2024a] (please see figure 15(b)). For the range of $\lambda_{eff}$ observed here, the variation in the $C_{P}$ and $C_{T}$ was expected to be marginal. The test cases are divided into 4 subgroups, primarily based on turbulence intensity. Group 1 has negligible $T_{i}$ while groups 2, 3 and 4 have low, mild and high $T_{i}$ respectively. Nevertheless, hard boundaries likely do not exist between the adjacent subgroups.
3 Effect of freestream turbulence on the dynamics of the wake
Instantaneous vorticity fields for 6 cases, case 1 (no grid case), 2a, 2c, 3b, 4b and 4d are shown in figure 6 for $\lambda_{\infty}=6$ (see also supplementary videos 1-6). These cases represent the wake at increasing turbulence intensities. The instantaneous vorticity fields show a rich variety of scales present in the flow. The individual tip vortices and the merged vortices are the most clearly observed for the no grid case (case 1) as highlighted in figure 6(a). Here, the tip vortices undergo merging at $x/D\approx 1.5$ similarly to the observation in our earlier study through a different set of experiments [Biswas and Buxton, 2024a]. As $T_{i}$ is increased, the individual tip vortices can still be observed (see figure 6(b,c)), but it becomes increasingly difficult to detect the merging process and the merged tip vortices. In general, as $T_{i}$ is increased, the tip vortices break down much closer to the turbine as reported in previous studies [Chamorro and Porté-Agel, 2009, Wu and Porté-Agel, 2012, Gambuzza and Ganapathisubramani, 2023]. For the no grid case, the wake edge towards the far wake (indicated by the black dash-dotted line) stays nearly flat. However, for the other cases the wake edge shows significantly larger undulation. This can be seen more clearly from the supplementary videos 1-6. We look at the important frequencies present at different locations in the wake to better understand these primitive observations.
Figure 7: Spectra at $x/D=0.5$ and $y/D=0.55$ for the cases (a) 1, (b) 2a, (c) 2c, and (d) 3b, (e) 4b and (f) 4d. The spectra with and without the turbine are shown in black and blue respectively. The shaded region shows the wake meandering frequency band ($0.15\leq St_{U_{b},D}\leq 0.4$).
Figure 8: Maximum power of the tip vortex related frequencies (a) $3f_{r}$ and (b) $f_{r}$ for the different FST cases.
3.1 Important frequencies